# Enabling Training-Free Semantic Communication Systems with Generative Diffusion Models

### v2

@@ Abstract
Index Terms:
Semantic communication, deep joint source-channel coding, diffusion models, image transmission.
@@ I Introduction
Semantic communication (SemCom) has emerged as a promising paradigm for the next generation of wireless communication systems and has attracted significant research interest in recent years. The key idea of SemCom is to understand the meaning of the transmitted data and transmit the most relevant information to the receiver, which can significantly reduce the amount of data transmitted over the channel and support downstream tasks at the receiver [1], such as autonomous driving, metaverse, and smart cities.
@@
Benefiting from recent advances in artificial intelligence (AI) technologies, various SemCom systems have been proposed. Specifically, the authors in [2] proposed a deep joint source-channel coding (DeepJSCC) system for wireless image transmission, where the semantic encoder and decoder are implemented using convolutional neural networks (CNNs) to extract the semantic information behind the original pixels and reconstruct the original image. DeepJSCC achieves superior reconstruction performance compared to conventional digital communication systems and effectively mitigates the cliff effect. Following this work, the authors in [3] modified the architecture of the semantic encoder and decoder by introducing the powerful transformer backbone, significantly improving transmission quality. In addition, a contrastive learning-based SemCom framework [4] was proposed to train the semantic encoder and decoder, which enhances the semantic consistency during transmission. However, due to the use of the autoencoder architectures and the discriminative AI paradigm, these approaches struggle to achieve high communication efficiency, and also require extensive training on large datasets and various channel conditions, which significantly limit their performance and flexibility in practical systems.
@@
Fortunately, the recent emergence of generative artificial intelligence (GenAI) offers new opportunities to overcome this limitation. By learning to capture the underlying data distribution rather than direct input-to-label mappings in discriminative AI [5], GenAI enables the generation of high-dimensional data (e.g., images or text) from low-dimensional vectors. This allows SemCom systems to transmit minimal data while enabling the receiver to reconstruct the original content through conditional sampling from the learned distribution [6]. For example, the authors in [7] proposed a generative JSCC framework in which a generative adversarial network (GAN) is integrated into the decoder to enhance reconstruction quality by leveraging the semantic priors learned by the generator. Moreover, more powerful generative diffusion models (GDMs)[8, 9] have been introduced into DeepJSCC-based frameworks[10, 11], where the degraded images reconstructed by DeepJSCC serve as conditional inputs to guide the diffusion sampling process, resulting in more realistic and perceptually faithful reconstructions. However, these approaches remain dependent on a well-trained DeepJSCC system.
@@
To eliminate the dependency on training, recent studies have explored training-free generative SemCom frameworks. Specifically, the authors in [12] introduced a channel-aware GAN inversion method for semantic encoding and employed the same GAN generator for decoding, thus avoiding any encoder–decoder training on the communication task. Meanwhile, the works in [13, 14] proposed to use a pretrained image captioner to extract textual descriptions from source images. These captions, along with auxiliary visual features such as edge maps or latent vectors, are transmitted to the receiver, where the caption serves as a conditioning prompt to guide a GDM in reconstructing the original image. However, these training-free approaches lack robustness to channel noise, as the transmitted semantic conditions, such as captions or latent features, are easily corrupted during transmission, leading to degraded reconstruction quality.
@@
To overcome these limitations, we propose a fully training-free generative SemCom framework that leverages publicly available pretrained GDMs. Specifically, we design a semantic encoding and decoding method based on the sampling and inversion processes of the denoising diffusion implicit model (DDIM) [9], which introduces a two-stage forward diffusion strategy split between the transmitter and receiver to enhance robustness against channel noise. In addition, we optimize the number of sampling steps at the receiver to match the total noise level introduced by the channel noise. We further provide a brief analysis of this design, how channel noise shifts the latent distribution away from the latent distribution of GDM in ideal conditions, and give insights into how the proposed system mitigates this mismatch. Simulations on the Kodak dataset are conducted to demonstrate the superiority of the proposed system across a wide range of perceptual and distortion metrics. In particular, our method achieves over 50% performance gains in Fréchet Inception distance (FID) compared to baselines when the SNR is below 5 dB.
@@ II-A System model of SemCom
In this paper, we consider a typical SemCom system for wireless image transmission, where the transmitter and receiver are equipped with a semantic encoder and decoder, respectively. For the input RGB image $\bm{x}\in\mathbb{R}^{3\times W\times C}$, where $H$ and $W$ denote the image height and width, respectively, the semantic encoder first extracts the semantic information and directly maps it into to a $k$ complex-dimensional channel input signal $\bm{z}\in\mathbb{C}^{k}$, given by
@@
where $\mathcal{E}_{\theta}(\cdot)$ is the semantic encoder with parameters $\theta$. To evaluate the communication efficiency, we define the bandwidth compression ratio as $\text{BCR}=\frac{k}{N}$, where $N=3\times H\times C$ denotes the source bandwidth. Then, the channel input $\bm{z}$ is transmitted over a noisy channel, which is modeled as
@@ II-B Generative Diffusion Models
GDMs are a class of generative models that learns to generate data by gradually denoising from a pure noise distribution. Specifically, a typical diffusion model consists of a forward diffusion process with no learnable parameters and a reverse denoising process with a learnable neural network[8]. Exemplifying the image generation task with latent diffusion, the forward diffusion process gradually adds Gaussian noise to the training data $\bm{z}$, which can be expressed as
$q(\bm{z}_{t}|\bm{x}_{t-1})=\mathcal{N}(\sqrt{\alpha_{t}}\bm{z}_{t-1},(1-\alpha_{t})\bm{I}),$ (4)
@@
$\bm{z}_{T_{F}}\sim\mathcal{N}(\sqrt{\alpha_{T_{F}}}\bm{z}_{0},(1-\bar{\alpha}_{T_{F}})\bm{I}),$ (5)
@@
$\mathcal{L}(\omega)=\mathbb{E}_{\bm{z}_{0},\bm{\epsilon},t}\left[||\bm{\epsilon}-\epsilon_{\omega}(\bm{z}_{t},t)||^{2}\right],$ (8)
where $\bm{\epsilon}$ is the ground-truth noise to be subtracted from the image at time step $t$. After training, the model can generate new images by sampling from the noise distribution $\bm{z}_{T}\sim\mathcal{N}(0,\bm{I})$ and iteratively applying the reverse denoising process in (6) for $T$ steps or performing jump sampling using the DDIM [9] to accelerate the generation without compromising quality, which can be expressed as (9) in the next page.
@@ III Proposed System
In this section, we introduce the proposed training-free generative SemCom system built upon the stable diffusion model, a DDIM-based latent diffusion model. We first present the key components of the proposed system, and then elaborate the details of the semantic encoding and decoding processes. Finally, we provide an analysis of the proposed system.
@@
III-A Overview
Following the stable diffusion model, our proposed generative SemCom system comprises a CLIP text encoder $\text{CLIP}(\cdot)$, an image encoder $\mathcal{E}_{\text{SD}}(\cdot)$, a conditional U-Net model $\epsilon_{\omega}(\cdot,\cdot,\cdot)$, and an image decoder $\mathcal{D}_{\text{SD}(\cdot)}$, as Fig. Fig. 1 illustrates. At the transmitter side, the semantic encoder consists of a CLIP text encoder, an image encoder, and a U-Net model, which is used to extract the semantic features from the input image and generate the channel input signal. At the receiver side, there is also the same CLIP text encoder, U-Net model as the transmitter side, and an image decoder, which is used to reconstruct the image from the received signal. We note that all the components are pre-trained and fixed during the transmission process to align with the training-free design.
@@ III-A Overview
III-B Semantic Encoding via DDIM Inversion
To extract the semantic information from the input image, we first use the image encoder to extract the latent feature, i.e.,  $\bm{z}_{0}=\mathcal{E}_{\mathrm{SD}}(\bm{x})$. Unlike the work in [13] that directly transmits the latent feature $\bm{z}_{0}$ as well as the extracted caption, and then adds random noise in the diffusion forward process at the receiver side, we make three key modifications to improve the system performance as follows.
@@ III-B Semantic Encoding via DDIM Inversion
    We propose to use the DDIM inversion to add $T_{F,1}$ steps of deterministic noise to the latent feature $\bm{z}_{0}$ before transmission. This process can be derived by reversing the DDIM sampling process in (9), and can be written as (10) at the top of this page. The term $\bm{c}$ is a text prompt, $\bm{s}=\text{CLIP}(\bm{c})$ is the corresponding embedding produced by the CLIP text encoder, and $\epsilon_{\omega}(\bm{z}_{t},t,\bm{s})$ is the predicted noise at time step $t$ by the U-Net model conditioned on $\bm{s}$. Thanks to the same U-Net model used in the forward and reverse processes, we can remove the noise more easily than the random noise added in [13].
@@
    Instead of extracting and transmitting the image caption, we use a predefined text prompt such as “High quality, High res, 8K,” which our experiments found to be more effective in improving reconstruction quality. We note that $T_{F,1}$ forward diffusion steps are performed at the transmitter, and the obtained latent $\bm{z}_{T_{F,1}}$ is normalized by a scaling factor $\gamma=1/\sqrt{\tfrac{1}{2k}||\bm{z}_{T_{F,1}}||_{2}^{2}}$ to satisfy the unit average power constraint [^1] The factor $1/2$ comes from the conversion from real-valued to complex-valued signals.. The normalized latent $\bm{z}=\gamma\bm{z}_{T_{F,1}}$is then mapped into complex-valued symbols for transmission.
III-C Semantic Decoding via DDIM Inversion and Sampling
@@ III-D Analysis
Original Refer to caption PSNR/LPIPS
BPG+LDPC Refer to caption 26.50/0.4415
DeepJSCC Refer to caption 28.16dB/0.4286
WITT Refer to caption 28.23dB/0.4078
SD_SemCom (Latent+Text) Refer to caption 20.67dB/0.4366
Proposed Refer to caption 24.67dB/0.2370
@@ IV-A Settings
In simulations, we implement the semantic encoder and decoder using the pretrained stable diffusion 1.5 model [^2] https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5, the BCR of the proposed system is set to $1/96$, and the SNR varies from 0 dB to 20 dB. During each transmission, the text prompt is set to “High quality, High res, 8K” for all images with a guidance scale of 6. We empirically set the total steps of the noise scheduler of stable diffusion is $50$, and the number of forward steps $T_{F,1}$ and $T_{F,2}$ are set to 5 and 5, respectively. The number of denoising steps $T_{B}$ is set to decrease from $18$ to $10$ as the SNR increases from 0 dB to 20 dB. We evaluate the performance under the widely used Kodak dataset [^3] https://r0k.us/graphics/kodak/ using various metrics, including PSNR, MS-SSIM, LPIPS, PIEAPP, DISTS, FID, and KID. We compare the proposed system with several baselines, including a traditional BPG compression followed by 5G LDPC for channel coding, an improved DeepJSCC system in [7], WITT[3], and the training-free semantic communication system SD_SemCom [13] with the same stable diffusion model.

### v3

@@ I Introduction
Semantic communication (SemCom) has emerged as a promising paradigm for the next generation of wireless communication systems and has attracted significant research interest in recent years. The key idea of SemCom is to understand the meaning of the transmitted data and transmit the most relevant information to the receiver, which can significantly reduce the amount of data transmitted over the channel and support downstream tasks at the receiver [Semantic1], such as autonomous driving, metaverse, and smart cities.
@@
Benefiting from recent advances in artificial intelligence (AI) technologies, various SemCom systems have been proposed. Specifically, the authors in [deepjscc] proposed a deep joint source-channel coding (DeepJSCC) system for wireless image transmission, where the semantic encoder and decoder are implemented using convolutional neural networks (CNNs) to extract the semantic information behind the original pixels and reconstruct the original image. DeepJSCC achieves superior reconstruction performance compared to conventional digital communication systems and effectively mitigates the cliff effect. Following this work, the authors in [SWINJSCC] modified the architecture of the semantic encoder and decoder by introducing the powerful transformer backbone, significantly improving transmission quality. In addition, a contrastive learning-based SemCom Framework [Shunpu_SemCom] was proposed to train the semantic encoder and decoder, which enhances semantic consistency during transmission. However, due to the use of the autoencoder architectures and the discriminative AI paradigm, these approaches struggle to achieve high communication efficiency, and also require extensive training on large datasets and various channel conditions, which significantly limit their performance and flexibility in practical systems.
@@
Fortunately, the recent emergence of generative artificial intelligence (GenAI) offers new opportunities to overcome this limitation. By learning to capture the underlying data distribution rather than direct input-to-label mappings in discriminative AI [AIGC], GenAI enables the generation of high-dimensional data (e.g., images or text) from low-dimensional vectors. This allows SemCom systems to transmit minimal data while enabling the receiver to reconstruct the original content through conditional sampling from the learned distribution [10614204]. For example, the authors in [GAN_JSCC] proposed a generative JSCC framework in which a generative adversarial network (GAN) is integrated into the decoder to enhance reconstruction quality by leveraging the semantic priors learned by the generator. Moreover, more powerful generative diffusion models (GDMs)[ho2020denoising, song2021denoising] have been introduced into DeepJSCC-based frameworks[Yilmaz-Other-2024, Chen-arXiveessIV-2025], where the degraded images reconstructed by DeepJSCC serve as conditional inputs to guide the diffusion sampling process, resulting in more realistic and perceptually faithful reconstructions. However, these approaches remain dependent on a well-trained DeepJSCC system.
@@
To eliminate the dependency on training, recent studies have explored training-free generative SemCom frameworks. Specifically, the authors in [tang2024evolving] introduced a channel-aware GAN inversion method for semantic encoding and employed the same GAN generator for decoding, thus avoiding any encoder–decoder training on the communication task. Meanwhile, the works in [Cicchetti_MLSP-2024, tang2024retrieval, Li_SemCom] proposed to use a pretrained image captioner to extract textual descriptions from source images. These captions, along with auxiliary visual features such as edge maps or latent vectors, are transmitted to the receiver, where the caption serves as a conditioning prompt to guide a GDM in reconstructing the original image. However, these training-free approaches lack robustness to channel noise, as the transmitted semantic conditions, such as captions or latent features, are easily corrupted during transmission, leading to degraded reconstruction quality.
@@
To overcome these limitations, we propose a fully training-free generative SemCom framework that leverages publicly available pretrained GDMs. Specifically, we design a semantic encoding and decoding method based on the sampling and inversion processes of the denoising diffusion implicit model (DDIM) [song2021denoising], which introduces a two-stage forward diffusion strategy split between the transmitter and receiver to enhance robustness against channel noise. In addition, we optimize the number of sampling steps at the receiver to match the total noise level introduced by the channel noise. We further provide a brief analysis of this design, how channel noise shifts the latent distribution away from the latent distribution of GDM in ideal conditions, and give insights into how the proposed system mitigates this mismatch. Simulations on the Kodak dataset are conducted to demonstrate the superiority of the proposed system across a wide range of perceptual and distortion metrics. In particular, our method achieves over 50% performance gains in Fréchet Inception distance (FID) compared to baselines when the SNR is below 5 dB.
@@ II-A System model of SemCom
In this paper, we consider a typical SemCom system for wireless image transmission, where the transmitter and receiver are equipped with a semantic encoder and decoder, respectively. For the input RGB image $\bm{x}\in\mathbb{R}^{3\times H\times W}$, where $H$ and $W$ denote the image height and width, respectively, the semantic encoder first extracts the semantic information and directly maps it into to a $k$ complex-dimensional channel input signal $\bm{z}\in\mathbb{C}^{k}$, given by
@@
where $\mathcal{E}_{\theta}(\cdot)$ is the semantic encoder with parameters $\theta$. To evaluate the communication efficiency, we define the bandwidth compression ratio as $\text{BCR}=\frac{k}{N}$, where $N=3\times H\times W$ denotes the source bandwidth. Then, the channel input $\bm{z}$ is transmitted over a noisy channel, which is modeled as
@@ II-B Generative Diffusion Models
GDMs are a class of generative models that learns to generate data by gradually denoising from a pure noise distribution. Specifically, a typical diffusion model consists of a forward diffusion process with no learnable parameters and a reverse denoising process with a learnable neural network[ho2020denoising]. Exemplifying the image generation task with latent diffusion, the forward diffusion process gradually adds Gaussian noise to the training data $\bm{z}$, which can be expressed as
$q(\bm{z}_{t}|\bm{z}_{t-1})=\mathcal{N}(\sqrt{\alpha_{t}}\bm{z}_{t-1},(1-\alpha_{t})\bm{I}),$ (4)
@@
$\bm{z}_{T_{F}}\sim\mathcal{N}(\sqrt{\bar{\alpha}_{T_{F}}}\bm{z}_{0},(1-\bar{\alpha}_{T_{F}})\bm{I}),$ (5)
@@
[…]
where $\bm{\epsilon}$ is the ground-truth noise to be subtracted from the image at time step $t$. After training, the model can generate new images by sampling from the noise distribution $\bm{z}_{T}\sim\mathcal{N}(0,\bm{I})$ and iteratively applying the reverse denoising process in (6) for $T$ steps or performing jump sampling using the DDIM [song2021denoising] to accelerate the generation without compromising quality, which can be expressed as (9) in the next page.
@@ III Proposed System
Following the stable diffusion model, our proposed generative SemCom system comprises a CLIP text encoder $\text{CLIP}(\cdot)$, an image encoder $\mathcal{E}_{\text{SD}}(\cdot)$, a conditional U-Net model $\epsilon_{\omega}(\cdot,\cdot,\cdot)$, and an image decoder $\mathcal{D}_{\text{SD}(\cdot)}$, as Fig. 1 illustrates. At the transmitter side, the semantic encoder consists of a CLIP text encoder, an image encoder, and a U-Net model, which is used to extract the semantic features from the input image and generate the channel input signal. At the receiver side, there is also the same CLIP text encoder, U-Net model as the transmitter side, and an image decoder, which is used to reconstruct the image from the received signal. We note that all the components are pre-trained and fixed during the transmission process to align with the training-free design.
@@
III-A Semantic Encoding via DDIM Inversion
To extract the semantic information from the input image, we first use the image encoder to extract the latent feature, i.e., $\bm{z}_{0}=\mathcal{E}_{\mathrm{SD}}(\bm{x})$. Unlike the work in [Cicchetti_MLSP-2024] that directly transmits the latent feature $\bm{z}_{0}$ as well as the extracted caption, and then adds random noise in the diffusion forward process at the receiver side, we make three key modifications to improve the system performance as follows.
@@ III-A Semantic Encoding via DDIM Inversion
    We propose to use the DDIM inversion to add $T_{F,1}$ steps of deterministic noise to the latent feature $\bm{z}_{0}$ before transmission. This process can be derived by reversing the DDIM sampling process in (9), and can be written as (10) at the top of this page. The term $\bm{c}$ is a text prompt, $\bm{s}=\text{CLIP}(\bm{c})$ is the corresponding embedding produced by the CLIP text encoder, and $\epsilon_{\omega}(\bm{z}_{t},t,\bm{s})$ is the predicted noise at time step $t$ by the U-Net model conditioned on $\bm{s}$. Thanks to the same U-Net model used in the forward and reverse processes, we can remove the noise more easily than the random noise added in [Cicchetti_MLSP-2024].
@@
[…]
III-B Semantic Decoding via DDIM Inversion and Sampling
@@ III-C Analysis
Original
PSNR/LPIPS
BPG+LDPC
26.50/0.4415
DeepJSCC
28.16dB/0.4286
WITT
28.23dB/0.4078
SD_SemCom (Latent+Text)
20.67dB/0.4366
Proposed
24.67dB/0.2370
@@ IV-A Settings
In simulations, we implement the semantic encoder and decoder using the pretrained stable diffusion 1.5 model [^2] https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5, the BCR of the proposed system is set to $1/96$, and the SNR varies from 0 dB to 20 dB. During each transmission, the text prompt is set to “High quality, High res, 8K" for all images with a guidance scale of 6. We empirically set the total steps of the noise scheduler of stable diffusion is $50$, and the number of forward steps $T_{F,1}$ and $T_{F,2}$ are set to 5 and 5, respectively. The number of denoising steps $T_{B}$ is set to decrease from $18$ to $10$ as the SNR increases from 0 dB to 20 dB. We evaluate the performance under the widely used Kodak dataset [^3] https://r0k.us/graphics/kodak/ using various metrics, including PSNR, MS-SSIM, LPIPS, PIEAPP, DISTS, FID, and KID. We compare the proposed system with several baselines, including a traditional BPG compression followed by 5G LDPC for channel coding, an improved DeepJSCC system in [GAN_JSCC], WITT[SWINJSCC], and the training-free semantic communication system SD_SemCom [Cicchetti_MLSP-2024] with the same stable diffusion model.